# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v1

@@ Abstract
We present a low-cost quadruped manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning (RL) purely in simulation. The system comprises 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions, and a low-level policy for quadruped movement and limb-control, 2) a progressive policy expansion approach for solving the long-horizon task together with a teacher-student framework for efficient high-level training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
   With budget-friendly but limited reliability and performance hardware, and just one wrist-mounted RGB camera, the entire system fully trained in simulation achieves high success rates for long horizon tasks involving search, move, grasp, and drop-into, with fluid sim-to-real transfer in a wide variety of indoor and outdoor scenes and lighting conditions.Extensive real-world evaluations show that on the long horizon mobile manipulation tasks, our system achieves good performance when transferred to real both in terms of task success rate and execution efficiency. Finally, we discuss the necessity of our sim-to-real techniques for legged mobile manipulation, and show their ablation performance.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy running in different scenes in the real world, with large scene variations, including ground properties, backgrounds, and distractors etc. These scenes are not available in simulation training. The annotations at the bottom are added for understanding task progress stages and are not part of the input to the system.
@@ I Introduction
Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing. As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environment – commonly referred to as mobile manipulators. Such systems typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
@@
Several promising directions exist in training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer. Each of the directions has its pros and cons.
Imitation learning benefits from learning directly from human experts teleoperating the hardware, thus, reasonable behavior and some successes can be achieved even with a small amount of training. However, it can be expensive to collect demonstrations to cover all the combinatorial variations of the scene, the task setup, and the hardware. Reinforcement learning (RL) can automate policy optimization so minimal human effort is needed to adapt to the environment and task variations. Furthermore, the resulting policy is often more efficient than human teleoperation. However, mastering long-horizon tasks in the real world still poses a significant challenge. Directly training an RL policy in the real world can be challenging due to hardware constraints and safety issues. Having to rely on real robots, training efficiency can be low, and scalability is often limited. Additionally, a model trained in one physical setup may not generalize to another with for example different background or floor texture.
RL in simulation solves the scalability and adaptability issues by running massively parallel simulations with randomized setups. The difficulty lies in the sim-to-real gap. Shortcomings of the design often only shows up in real world testing, at the very end of the development cycle. Making things worse, several iterations of modeling, design tweaking, policy training, and real testing are often needed to close up a particular sim-to-real gap. Given such a long and often expensive pipeline of tuning steps, RL in simulation can be slow to arrive at a working prototype, and can be expensive in terms of engineering costs and server time for simulation and training. Therefore, the hope is to develop general techniques to bridge the sim-to-real gaps, so that efforts spent on developing the training pipeline and the robotic system for some tasks can be reused for training future tasks. This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost robotic system to solve long-horizon, intricate quadruped manipulation tasks. Weak and often faulty hardware, the complexity of legged manipulation, and the long horizon are some of the factors that would compound and significantly expand the sim-to-real gap, making this a daunting task.
To the best of our knowledge, this is the first working solution of a legged manipulator fully trained with RL in simulation, for long horizon tasks, with low cost hardware.
We use a quadruped instead of a wheel base, for it can adapt to different terrains, but manipulation is also more challenging on legs, due to the unstable legged base.
@@
This results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue. To mitigate this, we utilize a hierarchical design consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also rely on a teacher-student training framework similar to prior work [29, 59].
@@
The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world. To tackle the difficulties of long-horizon task learning, we also incorporate task decomposition as a form of privileged information, where boundaries between subtasks are defined. The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state. We also condition the student visuomotor policy on language instructions, allowing us to naturally command the robot to achieve different tasks.
@@
We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation). We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer. Unlike prior work [29, 52], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning. SLIM relies on pure RL, and thus it does not require a single demonstration trajectory. Our contributions are as follows:
@@
    We conduct an extensive set of simulated and real-world experiments, achieving higher success rate and faster task completion than baselines. In the real world, we demonstrate robust task success across a wide range of indoor and outdoor environments.
@@ III-A Observations and Actions
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of all joint position readings.
@@ III-B Hierarchical Policy Structure
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, and has a much shorter time horizon.
@@ IV-B Progressive Policy Expansion
Figure 3: Teacher Network Structure. On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. In the case of actor-critic learning, both critic and actor networks have the same architecture, with the difference in that the actor network outputs an action distribution while the critic outputs a scalar value. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e. $\Pi[k]\equiv\pi^{k}$.
@@ V The Student: Policy Distillation Guided RL
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$
To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion (e.g., std. for Gaussian). This makes our distillation loss the best of two worlds: imitation learning and entropy regularization.
@@
When designing the student’s representation model, we specifically take visual sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the visual sim2real gap (Fig. 4). The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This potentially leads to a better generalization ability. While there might be many choices for the vision bottleneck, from discrete latent codes [19] to canonical images [24], we choose to use a pair of segmentation and depth maps. There are mainly two reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss. Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a visual bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [29, 52].
@@ VII Sim-to-Real Gap Reduction Techniques
When training a policy in simulation, we would like to ensure that the policy also performs well in the real world. In other words, we must minimize the “sim2real gaps” as much as possible. Generally, we can divide the sim2real gaps for image-conditioned policies into two distinct categories: the dynamics and visual gaps. Here, we outline essential techniques that proved invaluable in addressing both gaps. We stress that the incorporation of such techniques is crucial to real-world performance. For brevity, we discuss a subset of the techniques employed. In-depth details for these techniques can be found in Appendix C.
@@ VII-A Dynamics Gap Reduction
The dynamics gap is essentially caused by misaligned transition functions $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ between the simulated and real worlds. This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance. In fact, even minor covariate shifts can result in task failure if the task requires fine-grained high precision coordination, e.g., grasping a small object. Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real. In particular, we use the following key techniques for addressing the issue: 1) arm joint tracking error minimization, 2) domain randomization, and 3) object perturbations.
@@
We used PID controller for the robot arm. As opposed to PD controllers which will always have gravity-induced tracking errors, the integral term of PID controllers allow for more robust pose-agnostic tracking given that the joint delta targets are sufficiently small for particular operating frequency. This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification. More details can be found in Appendix C-A.
@@ VII-B Visual Gap Reduction
The second source of sim2real gap for image-conditioned policies is the visual gap. This gap is caused by an RGB distribution mismatch between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various visual sim2real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
@@
Inspired by SECANT [10], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim2real transfer. Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image. We choose not to introduce higher-level image randomization/generation [53] because it might impact image semantics.
@@ VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX-250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig. 7, serving as the robot’s sole RGB vision feed. The depth stream is not used in this work. Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around 12K USD. With the non-edu version, total cost is even lower (around 6K USD) and reflects more closely the cost of mass produced parts..
@@ VIII-B Baselines
We compare SLIM with a number of baselines as detailed below:
@@
We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e. RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. We have therefore excluded it from the subsequent real-world experiments.

### v2

@@ Abstract
We present a low-cost legged mobile manipulation system that solves long-horizon real-world tasks, trained by reinforcement learning purely in simulation. This system is made possible by 1) a hierarchical design of a high-level policy for visual-mobile manipulation following instructions and a low-level policy for quadruped movement and limb control, 2) a progressive exploration and learning approach that leverages privileged task decomposition information to train the teacher policy for long-horizon tasks, which will guide an imitation-based student policy for efficient training of the high-level visuomotor policy, and 3) a suite of techniques for minimizing sim-to-real gaps.
   In contrast to previous approaches that use high-end equipment, our system demonstrates effective performance with more accessible hardware – specifically, a Unitree Go1 quadruped, a WidowX250S arm, and a single wrist-mounted RGB camera – despite the increased challenges of sim-to-real transfer. When fully trained in simulation, a single policy autonomously solves long-horizon tasks such as search, move, grasp, and drop-into, achieving nearly 80% success. This performance is comparable to that of expert human teleoperation on the same tasks but operates in a more efficient way, at 1.5 times the speed of human expert. The sim-to-real transfer is fluid across diverse indoor and outdoor scenes under varying lighting conditions. Finally, we discuss the key techniques that enable the entire pipeline, including efficient RL training and sim-to-real, to work effectively for legged mobile manipulation, and present their ablation results.
Figure 1: SLIM in Real. Snapshots of the same SLIM policy deployed in diverse real-world scenes, featuring significant variations in ground types, backgrounds, distractors, and other environmental factors. These scenes are not available in simulation training. The annotations at the bottom are added for understanding task progress and are not part of the input to the system.
@@ I Introduction
Until recently, robots have primarily been limited to static, highly repetitive tasks, often in industrial manufacturing. As the demand for autonomy grows, the need emerges for robots to navigate and interact with their environments - commonly referred to as mobile manipulators. Such systems can typically handle primitive tasks, such as searching, grasping, and transporting objects, to achieve overarching long-horizon goals.
@@
There are several promising methodologies for training autonomous mobile manipulators: imitation learning from human demonstrations, reinforcement learning (RL) in the real world, and RL in simulation with sim-to-real transfer. Each of them has its pros and cons.
Imitation learning (IL) benefits from learning directly from human experts teleoperating the hardware, and thus reasonable behaviors and some success in a restricted domain can be achieved even with a handful of trajectories. However, diverse behavior data is required to improve the policy’s robustness even for a single task. This makes IL costly, especially if we further want to cover all the combinatorial variations of scenes and tasks for robot ability scaling.
Reinforcement learning (RL) can potentially generate an infinite amount of training data by allowing the robot to explore and interact with the environment. Furthermore, the resulting policy is often more optimal than a human teleoperating policy. However, directly training an RL policy in the real world can be challenging due to hardware maintenance, reward specification, and safety issues. Data scale is also largely limited by the difficulty of deploying a robot fleet for collecting data in parallel. As a result, mastering long-horizon tasks in the real world still poses a significant challenge.
RL in simulation could address the scaling, generalization, and safety issues by running massively parallel simulation with randomized scenes and tasks. But to successfully deploy the trained policy on a real robot, the challenge lies in sim-to-real gaps. Therefore, we attempt to develop effective and general techniques to bridge the gaps. This work focuses on training an end-to-end RL policy in simulation, and zero-shot deploying in the real world on a relatively low-cost quadruped based robot to solve long-horizon, intricate manipulation tasks. The complexity of accurate manipulation on legs, the variations of the physical environment, the scene and task setups, and the long task horizon are some of the factors that compound and increase the difficulty of RL training. On top of these factors, limited hardware reliability further exacerbates the sim-to-real gap, making this a daunting task.
@@
We train an end-to-end model that is able to achieve long-horizon tasks consisting of searching, grasping, transporting and dropping primitive tasks (Figure 1). For robot mobility, we choose a quadruped instead of a wheeled base, as it has the potential to adapt to many different types of terrains. The choice of a quadruped base results in a large 19 DOF (12 leg joints, 6 arm joints, and 1 gripper) system, exacerbating the already difficult long-horizon RL exploration issue. To mitigate this, we utilize a hierarchical teacher-student design, consisting of a high-level mobile manipulation policy and low-level quadruped controller, and also a teacher-student training framework [6] for the high-level policy. To the best of our knowledge, this is the first demonstration of a low-cost real-world quadruped manipulator fully trained with RL in simulation, for solving long-horizon tasks.
@@
The teacher policy has access to privileged information (e.g., structured and accurate object features) that is unavailable in the real world. To tackle the difficulties of long-horizon task learning, we also incorporate task decomposition as a form of privileged information, where boundaries between subtasks are defined. The student policy is then trained to distill the teacher policy while maximizing task rewards, but with access to only standard sensor data such as camera feed and proprioceptive state. We also condition the student visuomotor policy on language instructions, providing a natural way to the user for specifying the task to be executed by the robot.
@@
We name our robotic system SLIM (Sim-to-Real Legged Instructive Manipulation). We reiterate that SLIM is a complete robotic system with fluid sim-to-real transfer. Unlike prior work [30, 53, 59], SLIM learns all modules in simulation from scratch, without the need for any pretrained model or real-world finetuning. SLIM relies on pure RL, and thus it does not require a single demonstration trajectory. Our contributions are as follows:
@@
    We conduct an extensive set of experiments with real-world evaluations, achieving higher success rate and faster task completion than baselines. We also demonstrate robust task success across a wide range of indoor and outdoor environments in real-world.
@@ III-A Observations and Actions
    A temporal stack of proprioceptive state vectors of the robot with a shape of $(N,D)$, where $D$ is the dimensionality of the concatenation of proprioceptive reading of all joints. $\mathbf{s}_{\textrm{r}}\!=\![\mathbf{s}_{\textrm{leg}},\mathbf{s}_{\textrm{arm}}]$, where $\mathbf{s}_{\textrm{leg}}$ denotes the leg-related proprioceptive states and $s_{\textrm{arm}}$ the arm-related proprioceptive states. When used as the input to the low-level locomotion policy, we overload $\mathbf{s}_{\textrm{leg}}$ to denote the leg-related proprioceptive states stacked according to the low-level frame-stacking size.
@@ III-B Hierarchical Policy Structure
Note that with this decomposition, the low-level policy no longer observes task specific inputs $(\mathbf{l},\mathbf{o})$, thus does not need to handle high dimensional image and language inputs.
@@ IV-B Progressive Policy Expansion
Figure 3: Teacher Policy Network Structure. On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. In the case of actor-critic learning, both critic and actor networks have the same architecture, with the difference in that the actor network outputs an action distribution while the critic outputs a scalar value. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e. $\Pi[k]\equiv\pi^{k}$.
@@ V The Student: Policy Distillation Guided RL
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$ (5)
To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std. for Gaussian). This makes our distillation loss the best of two worlds: imitation learning and entropy regularization.
@@
When designing the student’s representation model, we specifically take vision sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the vision sim2real gap (Fig. 4). The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This potentially leads to a better generalization ability. While there might be many choices for the vision bottleneck, from discrete latent codes [20] to canonical images [25], we choose to use a pair of segmentation and depth maps. There are mainly two reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss. Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a vision bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [30, 53].
@@ VII Sim-to-Real Gap Reduction Techniques
sim-to-real is notoriously difficult. This work handles two kinds of sim-to-real gaps: the physical dynamics and the visual gaps. For brevity, we discuss a subset of the techniques which are crucial in addressing sim-to-real and for successful real-world deployment. More techniques can be found in Appendix D.
@@ VII-A Dynamics Gap Reduction
The dynamics gap is essentially caused by misaligned transition functions $P(\mathbf{s}^{t}|\mathbf{s}^{t-1},\mathbf{a}^{t-1})$ between the simulated and real worlds. This difference in dynamics will result in the sim-trained policy experiencing covariate shift when deployed in the real world, which if large enough, can destroy policy performance. In fact, even minor covariate shifts can result in task failure if the task requires fine motor control, e.g., grasping a small object. Therefore, to minimize the dynamics gap, we must minimize the impact of the covariate shift when transferring to real. In particular, we use the following key techniques for addressing the issue: 1) arm joint tracking error minimization, 2) domain randomization, and 3) object perturbations.
@@
We used PID controller for the robot arm. As opposed to PD controllers which will always have gravity-induced tracking errors, the integral term of PID controllers allow for more robust pose-agnostic tracking given that the joint delta targets are sufficiently small for particular operating frequency. This allows us to greatly reduce the dynamics sim-to-real gap while avoiding system identification. More details can be found in Appendix D-A.
@@
Stationary Manipulations.
We make the quadruped stand as still as possible when manipulating objects, i.e. during grasp or drop into. This avoids the large sim-to-real gap of a walking and shaking dog when accurate manipulation is needed. To achieve this, when training low level policies to follow locomotion speed commands, we also make the quadruped enter a stationary mode with a balanced stance when the speed command is close to zero. Without stationary manipulation, not only will success rates suffer, but also it becomes unsafe to deploy, because the gripper fingers can hit the ground, damaging fingers and arm motors.
@@ VII-B Visual Gap Reduction
The second source of sim-to-real gap for image-conditioned policies is the vision gap. This gap is caused by an RGB distribution mismatch between simulated and real world pixels. Since we do not assume knowing the target scenes in advance, we have to ensure that the perception model is able to handle a wide range of vision scenarios. Accordingly, we apply various vision sim-to-real gap reduction techniques to student training, in addition to the visual information bottleneck in Section V
@@
Inspired by SECANT [11], we apply pixel-level perturbations to every RGB image to increase the robustness of the perception model and the chance of its successful sim-to-real transfer. Specifically, we will randomly select a transformation from the list (Gaussian Noise, SaltPepper Noise, Speckle Noise, Random Brightness, Gaussian Blur) and apply it to each RGB image. We choose not to introduce higher-level image randomization/generation [54] because it might impact image semantics.
@@ VIII-A Physical Robotic System
For our robotic system, we use a Unitree Go1 with a top-mounted WidowX250S manipulator. An Intel RealSense D435 camera is attached to the WidowX’s wrist via a 3D-printed mount, as shown in Fig. 7, serving as the robot’s sole RGB vision feed. The depth stream is not used in this work. Additionally, custom 3D-printed, elongated parallel fingers are fitted on the WidowX gripper to extend its reach. The inner surface of the fingers are padded with a thin layer of foam, which act to increase grasping friction. Overall, our entire robot system is relatively low-cost [^2] The total cost of development version of the quadraped (Go1-edu) plus the arm is around \$12K USD. With the non-edu version, total cost is even lower ($\sim$6K USD) and reflects a bit more closely the cost of mass produced parts..
@@ VIII-B Baselines
We compare SLIM with a number of baselines below:
@@
  • Human Teleop: an expert human trained for half an hour in teleoperating the robot with locomotion and arm end-effector commands observing only the view of the wrist camera. While not a baseline per se, human teleoperation provides a valuable reference point. See Appendix B-D for more details.
We also tried another baseline No Distillation, which trains the visuomotor policy directly via RL without distillation from the teacher policy (i.e. RL for visuomotor policy learning from scratch). This baseline cannot learn to solve the task at all in simulation, because of the compounded difficulties of representation learning, behavior learning and long-horizon exploration. It is excluded from the tables.